OpenStudy (anonymous):

Salina currently has an account balance of $1,047.69. Her initial deposit on the account was $630 and it earned 3.9% simple interest. How long has Salina held the account?

OpenStudy (anonymous):

\[A = \frac{P \times r \times t}{100}\] A = 1047.69 P = 630 r = 3.9 Put these values and solve for t...

OpenStudy (anonymous):

I suck at fractions, i have